Suppose that the Market for Cigarette is facing the Demand function Q = 20 – 2P and Supply function Q = 10.5 + 0.5P: a) What is the effect on the Equilibrium Price and Quantity when Government imposes a 7% of tax as percent of equilibrium price on each unit of Cigarette produced? [5 marks] b) What is the price elasticity of demand at equilibrium after tax and comment on the answer? [5 marks]


1
Expert's answer
2021-04-14T11:21:55-0400

a)


"20-2p=10.5+0.5p"

"2.5p=9.5"

"p=3.8"

"p_t=1.07\times3.8=4.066"

"Q=11.868"

b)


"E=-2\times\frac{4.066}{11.868}=-0.68"

Since the government introduces an additional tax to reduce the consumption of cigarettes, we see that this measure is effective.